Give the center and radius of the cirle.
(x+.5)^2 + (y-6)^2=35

Answer by venugopalramana(3286)

Give the center and radius of the cirle.
(x+.5)^2 + (y-6)^2=35
STANDARD EQN.OF CIRCLE IS
(X-H)^2+(Y-K)^2=R^2....WHERE (H,K) IS THE CENTRE AND R=RADIUS...COMPARING
H=-0.5...K=6.....R=SQRT.(35)
SO CENTRE IS (-0.5,6) AND RADIUS =SQRT.(35)